# Fast Differentiable Modal Simulation of Non-linear Strings, Membranes, and Plates

### v1

@@ 2 Related Work
The simulation and modelling of vibrating structures such as strings, membranes, and plates has been extensively studied in acoustics and musical instrument research. Notably, the work of Bilbao [1] provides a comprehensive numerical treatment of such systems, primarily focusing on finite difference time-domain (FDTD) approaches. These methods discretise space and time, approximating derivatives using finite differences, and have been widely applied to the simulation of musical instruments and other acoustical systems. While FDTD methods are flexible and conceptually straightforward, they suffer from numerical dispersion and have restrictive stability conditions (e.g., CFL) that demand very small time steps. These constraints make high-frequency simulations or simulations involving large spatial domains computationally expensive. Recently, techniques like the Scalar Auxiliary Variable (SAV) method have also been introduced, primarily within FDTD frameworks, to improve the efficiency and stability of handling non-linear terms by avoiding implicit equations [2, 3].
Alternatively, modal approaches have been explored in different ways. Pseudospectral methods, which exploit Fourier decompositions, have been used for the simulation of the tension modulated string [4]. Another significant modal approach is the Functional Transformation Method (FTM), introduced by Trautmann and Rabenstein [5]. This technique relies on analytical spatial modes and the Laplace transform. By applying both transformations, the PDE becomes a set of transfer functions, which are then discretised and used in a numerical time integration scheme. However, analytical modal expansions exist only for specific geometries and boundary conditions. Subsequent works have focused on deriving analytical modes for more complex boundary conditions [6, 7]. Efficient tension-modulated modal models for strings and membranes have also been investigated in the works of Avanzini [8, 9, 10].
@@
In recent years, differentiable and neural-based approaches have emerged, primarily targeting rigid-body simulations [13, 14, 15]. Although most of these methods are primarily data-driven (rather than physics-based), they frequently leverage principles similar to classical modal expansions. In [16], non-linear modes are computed using gradient descent, and these modes are then evolved linearly, with a second neural network refining the dynamics to account for non-linear interactions. Similarly, Lee [17] proposes using analytical (or alternatively learnt) modes as a basis, modulating them with neural networks to model pitch glide and non-linear coupling effects. Such hybrid approaches highlight the potential benefits of integrating physics-based modal techniques with neural methods.
@@ 3 Background
For the modal approach, a suitable set of eigenfunctions (modal basis) must be found by solving the spatial eigenvalue problem:
$\Delta\Phi_{\mu}=\lambda_{\mu}\Phi_{\mu},$ (2)
@@ 3.1 Modal expansion of the non-linear term
For the von Kármán plate, both transverse modes $\Phi_{\mu}$ and in-plane modes $\Psi_{\mu}$ must be considered. The non-linear modal expansion then becomes:
@@
where $H_{q,r}^{n}$ and $C_{p,n}^{s}$ are third-order tensors that contain the coupling coefficients obtained from the projections between the modes:
@@ 3.2 Integration in time
$\omega_{\mu}^{2}$ $=\frac{D\lambda_{\mu}^{2}-T_{0}\lambda_{\mu}}{\rho},$ (13)
@@ 4.1 Implementation
The open source repository includes utilities for computing analytical modes and eigenvalues of strings, membranes, and plates for the clamped string and membrane cases. For plates, we provide a Python implementation of the approach described in [21] to compute non-linear coupling coefficients  [^1] https://vkgong.ensta-paris.fr. To support arbitrary boundary conditions, we use the magpie-python library  [^2] https://github.com/Nemus-Project/magpie-python together with a reimplementation for processing numerical modes based on  [^3] https://github.com/Nemus-Project/VKPlate.
@@ 4.2 Benchmarks
A key motivation for this work is to develop a high-performance simulation framework for non-linear models, particularly aimed at efficient large-scale dataset generation. To evaluate computational performance, we benchmark our Störmer-Verlet time integration approach (computationally equivalent to the discretised FTM method) for the von Kármán plate model implemented in JAX against three alternative implementations: a MATLAB version, an optimised C++ implementation using the Eigen library with BLAS support, and a JIT-compiled PyTorch implementation. This comparison is particularly challenging due to the tensor contraction (in Equation 9) required at each time step. As shown in Figure 2, while our JAX-based approach is moderately slower than the MATLAB and C++ implementations for smaller mode counts ($\lessapprox 50$), it demonstrates better efficiency as the number of modes increases. It also substantially outperforms the optimised PyTorch implementation running on the GPU. Notably, at approximately 100 modes, our implementation is the only one capable of achieving roughly twice real-time performance for a one-second simulation at a 44,100 Hz sampling rate.
@@ 4.3.1 Loss functions and setup
where $x$ and $\hat{x}$ represent the time-domain target and predicted signals, respectively, with $Y=\text{STFT}(x)$ and $\hat{Y}=\text{STFT}(\hat{x})$ denoting their corresponding STFT magnitude spectrograms. The terms $Y_{i}$ and $\hat{Y}_{i}$ refer to individual (normalised) time frames of the spectrograms used in the spectral optimal transport loss $\mathcal{L}_{\text{sot}}$ via the Wasserstein-1 distance $\mathcal{W}_{1}$. The constant $\epsilon$ ensures numerical stability in the logarithmic term. The weights $\alpha$, $\beta$, and $\eta$ are calibrated for each experiment to balance the contribution of each loss component. For linear cases, we also compute a slight variation of these losses where the STFT is replaced by the sampled magnitude frequency response, which provides a more direct measure of the systems transfer function (available in this specific case). Although $\mathcal{L}_{\text{sot}}$ helps avoid poor local minima, the overall spectral loss landscape remains non-convex, as illustrated in Figure 3, even for a single parameter.
@@ 4.3.2 String
In the first approach, and in particular for the real-world cases, we extract the spectral envelope from the target response using high-order Linear Predictive Coding (LPC), though alternative methods could be employed. We then sample the LPC coefficients in frequency using the Bark scale. On the model side, we compute the transfer function of the linear system using eq. 15, based on a set of parameters to be optimised. The predicted response is sampled in the same way as the LPC envelope. We then compute the loss using the functions defined in eq. 24, replacing the STFT magnitude with the Bark-sampled magnitude response. This provides a simple and efficient route for inverse modelling in the linear case, avoiding time-domain simulation. Figure 4a shows the semi-log plot of the fit to a real plucked string.
@@
We fit the parameters of a single recording of a real plucked string from the IDMT-SMT-Guitar dataset [23]. For this experiment, we assume the string is clamped at both ends and that nonlinear effects (e.g. pitch glide) are minimal.
While optimisation converges more quickly in the synthetic case, the real-world scenario presents greater challenges. This increased complexity arises from the need to account for not only the PDE parameters but also the external force excitation, the acoustic transfer function, measurement noise, and other effects. We also assume the string clamped at both ends.
To address the unknown initial conditions, we extend the model by optimising an additional parameter $b_{\mu,2}$ (in Equation 15) for each mode. Furthermore, instead of projecting the solution back to the physical domain using analytical mode shapes, we simply optimise a set of abstract weights for each mode. These weights account for effects such as the acoustic transfer and microphone placement, even if they are not physically interpretable.
@@ 4.3.3 Plate
In the synthetic non-linear case, we simulate the displacement of the plate at a fixed point in the domain. We then optimise the parameters by propagating gradients through the time integrator to match this response (i.e., using BPTT). This type of optimisation is more sensitive to parameter initialisation of parameters and is prone to getting trapped in local minima. Even when optimising a single parameter, such as $\hat{D}$, the loss landscape exhibits many local minima (Figure 3). This behaviour is also consistent with previously observed challenges in the optimisation of frequencies [turian_im_2020]. In addition, care must be taken to ensure that all parameters remain within numerically stable ranges. In particular, the damping coefficients $\gamma_{\mu}$ must be constrained to ensure all poles lie in the left half of the Laplace domain to avoid numerical instability. To address the challenge of optimising the bending stiffness and avoiding poor local minima, we adjust the weighting of $\mathcal{L}_{\text{total}}$ to emphasise the spectral optimal transport and spectral convergence losses. Since the optimisation using automatic differentiation runs is fast (for approximately 0.1 seconds at 44,100 Hz), we can efficiently explore multiple random initialisations in parallel [25].
@@ 5 Conclusion
Extensive benchmarks demonstrate that our implementation scales efficiently with an increasing number of modes, providing substantial performance gains compared to existing MATLAB, optimised C++, and GPU-based PyTorch (JIT) implementations. Furthermore, it supports effective gradient-based optimisation of physical parameters from both synthetic simulations and real-world measurements. We successfully recovered parameters such as tension, bending stiffness, damping, and geometry, including the coupling tensors of the non-linear von Kármán plate model.

### v2

@@ 2 Related Work
The simulation and modelling of strings, membranes, and plates has been extensively studied in acoustics and musical instrument research. Notably, the work of Bilbao [1] provides a comprehensive numerical treatment of such systems, primarily focusing on finite difference time-domain (FDTD) approaches. These methods discretise space and time, approximating derivatives using finite differences, and have been widely applied to the simulation of musical instruments and other acoustical systems. While FDTD methods are flexible and conceptually straightforward, they suffer from numerical dispersion and have restrictive stability conditions (e.g., CFL) that demand very small time steps. These constraints make high-frequency simulations or simulations involving large spatial domains computationally expensive. Recently, techniques like the Scalar Auxiliary Variable (SAV) method have also been introduced, primarily within FDTD frameworks, to improve the efficiency and stability of handling non-linear terms by avoiding implicit equations [2, 3].
Alternatively, modal approaches have been explored in different ways. Pseudospectral methods, which exploit Fourier decompositions, have been used for the simulation of the tension modulated string [4]. Another significant modal approach is the Functional Transformation Method (FTM), introduced by Trautmann and Rabenstein [5]. This technique relies on analytical spatial modes and the Laplace transform. By applying both transformations, the PDE becomes a set of transfer functions, which are then discretised and used in a numerical time integration scheme. However, analytical modal expansions exist only for specific geometries and boundary conditions. Subsequent works have focused on deriving analytical modes for more complex boundary conditions [6]. In parallel, efficient tension-modulated modal models for strings and membranes have also been developed in the works of Avanzini and Marogna [7, 8, 9].
@@
In recent years, differentiable and neural-based approaches have emerged, primarily targeting rigid-body simulations [12, 13, 14]. Although most of these methods are primarily data-driven (rather than physics-based), they often leverage principles similar to classical modal expansions. In [15], non-linear modes are computed using gradient descent, and these modes are then evolved linearly, with a second neural network refining the dynamics to account for non-linear interactions. Similarly, Lee [16] proposes using analytical (or alternatively learnt) modes as a basis, modulating them with neural networks to model pitch glide and non-linear coupling effects. Such hybrid approaches highlight the potential benefits of integrating physics-based modal techniques with neural methods.
@@ 3 Background
For the modal approach, a suitable set of eigenfunctions must be found by solving the spatial eigenvalue problem:
$\Delta\Phi_{\mu}=-\lambda_{\mu}\Phi_{\mu},$ (2)
@@ 3.1 Modal expansion of the non-linear term
For the von Kármán plate, both the transverse modes $\Phi_{\mu}$ and in-plane modes $\Psi_{\mu}$ must be considered. The non-linear modal expansion then becomes:
@@
where $H_{q,r}^{n}$ and $C_{p,n}^{s}$ are third-order tensors with the coupling coefficients obtained from the projections between the modes:
@@ 3.2 Integration in time
$\omega_{\mu}^{2}$ $=\frac{D\lambda_{\mu}^{2}+T_{0}\lambda_{\mu}}{\rho},$ (13)
@@ 4.1 Implementation
The open source repository includes utilities for computing analytical modes and eigenvalues of fixed strings, membranes and plates. For plates, we provide a Python implementation of the approach described in [11] to compute non-linear coupling coefficients [^1] https://vkgong.ensta-paris.fr. To support arbitrary boundary conditions, we use the magpie-python library [^2] https://github.com/Nemus-Project/magpie-python together with a reimplementation for processing numerical modes based on [^3] https://github.com/Nemus-Project/VKPlate.
@@ 4.2 Benchmarks
A key motivation for this work is to develop a high-performance simulation framework for non-linear models, particularly aimed at efficient large-scale dataset generation. To evaluate computational performance, we benchmark our Störmer-Verlet time integration approach (computationally equivalent to the discretised FTM method) for the von Kármán plate model implemented in JAX against three alternative implementations: a MATLAB version, an optimised C++ implementation using the Eigen library with BLAS support, and a JIT-compiled PyTorch implementation. This comparison is particularly challenging due to the tensor contraction (in Equation 9) required at each time step. As shown in Figure 2, while our JAX-based approach is moderately slower than the MATLAB and C++ implementations for smaller mode counts ($\lessapprox 50$), it demonstrates better efficiency as the number of modes increases. It also substantially outperforms the optimised PyTorch implementation running on the GPU. Notably, for the von Kármán plate model with approximately 100 modes, our implementation is the only one capable of achieving roughly twice real-time performance for a one-second simulation at a 44,100 Hz sampling rate.
@@ 4.3.1 Loss functions and setup
where $x$ and $\hat{x}$ represent the time-domain target and predicted signals, respectively, with $Y=|\text{STFT}(x)|$ and $\hat{Y}=|\text{STFT}(\hat{x})|$ denoting their corresponding STFT magnitude spectrograms. The terms $Y_{i}$ and $\hat{Y}_{i}$ refer to individual (normalised) time frames of the spectrograms used in the spectral optimal transport loss $\mathcal{L}_{\text{sot}}$ [20] via the Wasserstein-1 distance $\mathcal{W}_{1}$. The constant $\epsilon$ ensures numerical stability in the logarithmic term. The weights $\alpha$, $\beta$, and $\eta$ are calibrated for each experiment to balance the contribution of each loss component. For linear cases, we also compute a slight variation of these losses where the STFT is replaced by the sampled magnitude frequency response, which provides a more direct measure of the systems transfer function (available in this specific case). Although $\mathcal{L}_{\text{sot}}$ helps avoid poor local minima, the overall spectral loss landscape remains non-convex, as illustrated in Figure 3, even for a single parameter.
@@ 4.3.2 String
In the first approach, and in particular for the real-world scenarios, we extract the spectral envelope from the target response using high-order Linear Predictive Coding (LPC), though alternative methods could be employed. We then sample the LPC coefficients in frequency using the Bark scale. On the model side, we compute the transfer function of the linear system using eq. 15, based on a set of parameters to be optimised. The predicted response is sampled in the same way as the LPC envelope. We then compute the loss using the functions defined in eq. 24, replacing the STFT magnitude with the Bark-sampled magnitude response. This provides a simple and efficient route for inverse modelling in the linear case, avoiding time-domain simulation. Figure 4a shows the semi-log plot of the fit to a real plucked string.
@@
We fit the parameters of a single recording of a real plucked string from the IDMT-SMT-Guitar dataset [22]. For this experiment, we assume the string is fixed at both ends and that nonlinear effects (e.g. pitch glide) are minimal.
While optimisation converges more quickly in the synthetic case, the real-world scenario presents greater challenges. This increased complexity arises from the need to account for not only the PDE parameters but also the external force excitation, the acoustic transfer function, measurement noise, and other effects.
We extend the model by optimising an additional parameter $b_{\mu,2}$ (in Equation 15) for each mode, to address the unknown initial conditions. Furthermore, instead of projecting the solution back to the physical domain using analytical mode shapes, we simply optimise a set of abstract weights for each mode. These weights account for effects such as the acoustic transfer and microphone placement, even if they are not physically interpretable.
@@ 4.3.3 Plate
In the synthetic non-linear case, we simulate the displacement of the plate at a fixed point in the domain. We then optimise the parameters by propagating gradients through the time integrator to match this response (i.e., using BPTT). This type of optimisation is more sensitive to parameter initialisation of parameters and is prone to getting trapped in local minima. Even when optimising a single parameter, such as $\hat{D}$, the loss landscape exhibits many local minima (Figure 3). This behaviour is also consistent with previously observed challenges in the optimisation of frequencies [24]. In addition, care must be taken to ensure that all parameters remain within numerically stable ranges. In particular, the damping coefficients $\gamma_{\mu}$ must be constrained to ensure all poles lie in the left half of the Laplace domain to avoid numerical instability. To address the challenge of optimising the bending stiffness and avoiding poor local minima, we adjust the weighting of $\mathcal{L}_{\text{total}}$ to emphasise the spectral optimal transport and spectral convergence losses. Since the optimisation using automatic differentiation runs is fast (for approximately 0.1 seconds at 44,100 Hz), we can efficiently explore multiple random initialisations in parallel [25].
@@ 5 Conclusion
The benchmarks demonstrate that our implementation scales efficiently with an increasing number of modes, providing substantial performance gains compared to existing MATLAB, optimised C++, and GPU-based PyTorch (JIT) implementations. Furthermore, it supports effective gradient-based optimisation of physical parameters from both synthetic simulations and real-world measurements. We successfully recovered parameters such as tension, bending stiffness, damping, and geometry, including the coupling tensors of the non-linear von Kármán plate model.